I really need help with this fast!
The variable z varies jointly with x and y and inversely with the square of t. Also, z=5 when x=2, y=-3, and t=1/2. Write a function such that f(y)=t when x=-12/(5z)^2 and z=2/5.

f(y)=________
please explain and thanks in advance.

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39799)

Already answered. You need to be precise.

Answer by stanbon(75887)

The variable z varies jointly with x and y and inversely with the square of t.
z = k*x*y/t^2
------------
Also, z=5 when x=2, y=-3, and t=1/2.
Use these value to solve for "k"::
5 = k*2*-3/(1/2)^2
---
5 = k*-6/4
-6k = 20
-k = 10/3
k = 10/3
-------------------
Equation::
z = (10/3)*x*y/t^2
===============================
Write a function such that f(y)=t when x=-12/(5z)^2 and z=2/5.
---------
(2/5) = (10/3)*(12/25z^2)*y/t^2
----------------
Solve for "y"::::
y = (2/5)t^2/[(8/5z^2]
y = (1/4)(t/z)^2
Cheers,
Stan H.




f(y)=________
======
Cheers,
Stan H.
-----------------